# PSCon: Toward Conversational Product Search

### v1

@@ 5.2.5. Item Ranking (T5)
Similar to system action prediction and question selection, the sequence of item ranking is defined as $S_{T5}=[[T5],C,U,D]$. We first obtain the hidden representation of [T5] from the transformer base model. Then, similar to (T4), we apply a softmax function through a feedforward network with ReLU activation in between (Equation 3) to calculate the distribution probability over candidate items. Again, we model it as a ranking problem, one can also model it as a binary classification problem to predict whether each candidate item is selected or not.
@@ 5.2.6. Response Generation (T6)
We use a standard Transformer decoder to conduct response generation. The transformer decoder outputs a representation at each decoding time step. For response generation, we put the predicted system action (T3) at the beginning to indicate the type of response. We use the representations of T1 (user intent detection), T4 (question selection), and T5 (item ranking) as the representations of input sequences and fuse them with max pooling to get the final hidden representation. Then a linear classifier with softmax is applied on top to predict the probability of the token at each time step. In addition, we deploy the copy mechanism (Vinyals et al., 2015), which is commonly used and helpful for generating informative utterances (Gu et al., 2016), to copy tokens in the conversation history $C$, and product-related documents $D$.
@@ 6.1.1. Evaluation Metrics
The used evaluation metrics differ for different subtasks. For keyword extraction (T2) and response generation (T6), we use BLEU-1 (Papineni et al., 2002) and ROUGE-L (Lin, 2004) as they are supposed to evaluate the quality of texts following previous work (Zhao et al., 2023). For user intent detection (T1), and system action prediction (T3), we use Precision, Recall, and F1 as evaluation metrics as they are supposed to evaluate the quality of classification. For question selection (T4), and item ranking (T5), we use MAP and nDCG as evaluation metrics to evaluate the ranking quality of questions and items. nDCG is calculated by the top 100 items like other papers  (Van Gysel et al., 2016).
@@ 6.1.2. Implementation Details
We run the experiments in the NVIDIA RTX A6000*8 GPU server. We use Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.001 to train our model. The word embedding size and hidden size are set to 512. For Transformer, the number of layers for the encoder and decoder is set to 4 and 2, respectively. We set the Transformer head number = 8, and the global norm clip of gradients = 1 for stable training. We use 5-fold cross-validation to evaluate our model. In each fold, the dataset is split into training, valid, and testing set by the ratio of 6:2:2. In the following experiments, we omit the results for the English dataset and only report results for the Chinese dataset due to similar trends and limited space.
@@ 6.2.1. Performance on Different Tasks
We report the results of our methods on the collected Chinese dataset on all the six defined subtasks, as shown in Table 2. As there is a lack of existing conversational product search baselines that can be applied to all six subtasks, we include a set of typical and well-known classification baselines, including naive Bayes classifier, logistic regression classifier, random forest classifier, and Support Vector Machine (SVM) classifier, on the subtasks modeled by classification problems (i.e., T1-T3), while including a typical ranking baseline, Multilayer Perceptron (MLP), on the subtasks modeled by ranking problems (i.e., T4). Also, large language models (LLMs) can be employed as a solution which is a supplement to our proposed model. Therefore, we include the advanced pre-trained LLMs Llama 2 and GPT-4, by following the same setting in He et al. (2023), on our PSCon dataset.101010 The prompts used for pre-trained LLMs: “suppose your task is {subtask}. I will give you some contextual information. Using the contextual information as input data, predict the {subtask output}. The contextual information is provided as follows: {subtask input}.” In the prompt, the {subtask} is user intent detection, keyword extraction, system action prediction, question selection, item ranking, and response generation for T1, T2, T3, T4, T5, and T6, respectively. The {subtask output} is user intent, keywords, system action, ranked list of questions, ranked lists of items, and response for T1, T2, T3, T4, T5, and T6, respectively. The {subtask input} is shown in Section 4 for each subtask. The results aim to provide practitioners with a better perception of what are the advantages and challenges of using the dataset.
@@
From the results, we observe that random forest and SVM perform the best among the four typical classification baselines for classification subtasks (T1-T3). Compared with typical classification baselines, GPT-4 and Llama 2 achieve relatively better performance in T2 and T6 while worse performance in T4 and T5. This might be because pre-trained LLMs like GPT-4 are good at natural language processing tasks (e.g., T2 and T6) but may not be good at question selection and item selection over a large number of candidate questions and items without training on the downstream datasets. With fine-tuning on the PSCon dataset, the LLM Llama 2 greatly improves its performance on the six subtasks, by achieving higher performance than GPT-4 and Llama 2 without fine-tuning.
@@
Moreover, we observe that the PSCon dataset is valid for examining all subtasks and the proposed model is effective for the six subtasks. Our model, although a much smaller parameter size than LLMs like GPT-4 and Llama 2, achieving the best performance on most subtasks, constitutes a straightforward baseline for future researchers to use. The model is designed as a unified model to model the six subtasks in a unified framework. But note that the model is not the focus of the work. Overall, the subtasks of user intent detection (T1), action prediction (T3), question selection (T4), item ranking (T5), and response generation (T6) achieve relatively high performance, indicating the effectiveness of our PSCon dataset on these subtasks. Using the conversation history for user intent detection, action prediction, question selection, item ranking, and response generation is beneficial. We observe higher effectiveness in terms of Precision for system action prediction, suggesting the need for further research in this area to improve recall. We also observe that the item ranking is challenging and far from perfect. This might be because of the large number of candidate items, which is a known problem in information filtering and recommender systems, and calls for the need to employ more complex models that learn from user-item interactions, such as collaborative filtering (He et al., 2017). Finally, we think there is a lot of room for improvement in response generation, given the low performance of GPT-4 and our model in terms of BLEU and ROUGE, indicating that CPS is a promising direction but there is still a long way to go. To further validate the response generation, we add a case study shown in Figure 6. From the response generation example, we observe that, while GPT-4 generates general responses omitting product recommendations, our model generates reasonable responses with specific product recommendations, which are similar to the ground truth. However, the responses generated by our model do not contain the same phrase with the ground truth although their meaning is the same, leading to the relatively low BLEU and ROUGE.
@@ 6.2.3. Performance on System Action Prediction
The detailed performance of our model in terms of each system action, namely, ‘Chitchat’, ‘Clarify’, ‘Recommend’, and ‘No-answer’, are shown in Table 4. We observe that the performance changes for different system actions. Specifically, the model achieves the highest F1 when the system actions are ‘Recommend’ and ‘Clarify’. This might be because ‘Recommend’ and ‘Clarify’ actions are less ambiguous and can be easily classified.
@@ 7. Conclusion and Future Work
To collect the data, two workers need to be paired as participants to generate one conversation. Each conversation usually involves multiple searches and scanning for products which costs much effort of the participants. This makes the data collection challenging and expensive, leading to generating a limited volume of data like existing work (Trippas et al., 2020; Thomas et al., 2017; Dalton et al., 2020; Bernard and Balog, 2023). We understand that, although the data contains more than one thousand conversations, the data is still limited to training a large model in an end-to-end manner. To this end, we utilize other conversational datasets to pre-train our model to relax this limitation and the experimental results of our proposed model demonstrate its effectiveness. Nevertheless, we plan to maintain the dataset by expanding it with more training data from a large number of users in the future. As we present a system and data collection protocol in the paper, one can easily collect more conversations by using our system and protocol. Moreover, more than one thousand conversations enable researchers to use the dataset with LLMs (Mao et al., 2024) and unlock its capabilities, e.g., one can employ LLM’s few-shot learning by carefully curating domain-specific examples and prompts from the dataset; one can also use LLM’s knowledge injection by integrating our conversations as demonstrations or leveraging our knowledge graphs through retrieval-augmented generation (RAG) to enhance its reasoning and contextual understanding.
@@
In this paper, we experiment with our PSCon dataset to offer an overview of the difficulty and potential of our PSCon for targeting different tasks. Nevertheless, we do not experiment with cross-lingual or cross-market models (Zhu et al., 2022; Bonab et al., 2021) on our datasets. As we collect two datasets from two markets with two languages, there are several items from the international e-commerce stores that are the same across the two datasets. One can use the conversation in one market with a certain language to retrieve the target products in the other market with a different language. The information and knowledge of items (e.g., the conversation, product metadata, and product reviews) can transfer from one dataset to another dataset, which can be used for training domain-adaptation product search models, especially when one dataset provides a limited amount of training data. Also, the item interactions from conversations in one dataset can be learned and used as collaborative signals (e.g., users like Roccat Kain 120 AIMO mouse may also like Roccat Kone XP Air mouse, as stated in our PSCon dataset) in the other dataset for improving cross-domain (i.e., cross-lingual or cross-market) product recommendations. We plan to continuously maintain the dataset by expanding it with more languages/markets and train cross-domain product search models (e.g., (Bonab et al., 2021) and (Zhu et al., 2022)) on our datasets in our future work.
Although we have designed six subtasks for CPS and proposed a unified Transformer model to model these six subtasks, the proposed model is straightforward and simple. It constitutes a straightforward baseline for future researchers to use. Nevertheless, it is worth mentioning again that CPS is a promising direction and there is still a long way to go to improve the performance of each subtask. Our work is a preliminary study and may not represent a perfect setting of CPS, but we view it as a stepping stone toward more advanced CPS.

### v2

@@ 5.2.5. Item Ranking (T5)
Similar to system action prediction and question selection, the sequence of item ranking is defined as $S_{T5}=[[T5],C,U,D]$. We first obtain the hidden representation of [T5] from the transformer base model. Then, similar to (T4), we apply a softmax function through a feedforward network with ReLU activation in between (Equation 3) to calculate the distribution probability over candidate items. Again, we model it as a ranking problem.
@@ 5.2.6. Response Generation (T6)
We use a standard Transformer decoder to conduct response generation. The transformer decoder outputs a representation at each decoding time step. For response generation, we put the predicted system action (T3) at the beginning to indicate the type of response. We use the representations of (T1), (T4), and (T5) as the representations of input sequences and fuse them with max pooling to get the final hidden representation. Then, a linear classifier with softmax is applied on top to predict the probability of the token at each time step. In addition, we deploy the copy mechanism (Vinyals et al., 2015), which is commonly used and helpful for generating informative utterances (Gu et al., 2016), to copy tokens in the conversation history $C$, and product-related documents $D$.
@@ 6.1.1. Evaluation Metrics
The evaluation metrics used differ for different subtasks. For keyword extraction (T2) and response generation (T6), we use BLEU-1 (Papineni et al., 2002) and ROUGE-L (Lin, 2004) as they are supposed to evaluate the quality of texts following previous work (Zhao et al., 2023). For user intent detection (T1), and system action prediction (T3), we use Precision, Recall, and F1 as evaluation metrics as they are supposed to evaluate the quality of classification. For question selection (T4), and item ranking (T5), we use MAP and nDCG as evaluation metrics to evaluate the ranking quality of questions and items.
@@ 6.1.2. Implementation Details
We run the experiments on the NVIDIA RTX A6000*8 GPU server. We use Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.001 to train our model. The word embedding size and hidden size are set to 512. For Transformer, the number of layers for the encoder and decoder is set to 4 and 2, respectively. We set the Transformer head number = 8, and the global norm clip of gradients = 1 for stable training. We use 5-fold cross-validation to evaluate our model. In each fold, the dataset is split into training, validation, and testing sets by the ratio of 6:2:2. In the following experiments, we omit the results for the English dataset and only report results for the Chinese dataset due to similar trends and limited space.
@@ 6.2.1. Performance on Different Tasks
We report the results of our methods on the collected Chinese dataset on all the six defined subtasks, as shown in Table 2. As there is a lack of existing conversational product search baselines that can be applied to all six subtasks, we include a set of typical and well-known classification baselines, including naive Bayes classifier, logistic regression classifier, random forest classifier, and Support Vector Machine (SVM) classifier, on the subtasks modeled by classification problems (i.e., T1-T3), while including a typical ranking baseline, Multilayer Perceptron (MLP), on the subtasks modeled by ranking problems (i.e., T4). Also, large language models (LLMs) can be employed as a solution which is a supplement to our proposed model. Therefore, we include the advanced pre-trained LLMs Llama 2 (Llama-2-7B) and GPT-3.5, by following the same setting in He et al. (2023), on our PSCon dataset.888 The prompts used for pre-trained LLMs: “suppose your task is {subtask}. I will give you some contextual information. Using the contextual information as input data, predict the {subtask output}. The contextual information is provided as follows: {subtask input}.” In the prompt, the {subtask} is user intent detection, keyword extraction, system action prediction, question selection, item ranking, and response generation for T1, T2, T3, T4, T5, and T6, respectively. The {subtask output} is user intent, keywords, system action, ranked list of questions, ranked lists of items, and response for T1, T2, T3, T4, T5, and T6, respectively. The {subtask input} is shown in Section 4 for each subtask. The results aim to provide practitioners with a better perception of what are the advantages and challenges of using the dataset.
@@
From the results, we observe that random forest and SVM perform the best among the four typical classification baselines for classification subtasks (T1-T3). Compared with typical classification baselines, GPT-3.5 and Llama 2 achieve relatively better performance in T2 and T6 while worse performance in T4 and T5. This might be because pre-trained LLMs like GPT-3.5 are good at natural language processing tasks (e.g., T2 and T6) but may not be good at question selection and item selection over a large number of candidate questions and items without training on the downstream datasets. With fine-tuning on the PSCon dataset, the LLM Llama 2 greatly improves its performance on the six subtasks, by achieving higher performance than GPT-3.5 and Llama 2 without fine-tuning.
@@
Moreover, we observe that the PSCon dataset is valid for examining all subtasks and the proposed model is effective for the six subtasks. Our model, although a much smaller parameter size than LLMs like GPT-3.5 and Llama 2, achieving the best performance on most subtasks, constitutes a straightforward baseline for future researchers to use. The model is designed as a unified model to model the six subtasks in a unified framework. But note that the model is not the focus of the work. Overall, the subtasks of user intent detection (T1), action prediction (T3), question selection (T4), item ranking (T5), and response generation (T6) achieve relatively high performance, indicating the effectiveness of our PSCon dataset on these subtasks. Using the conversation history for user intent detection, action prediction, question selection, item ranking, and response generation is beneficial. We observe higher effectiveness in terms of Precision for system action prediction, suggesting the need for further research in this area to improve recall. We also observe that the item ranking is challenging and far from perfect. This might be because of the large number of candidate items, which is a known problem in information filtering and recommender systems, and calls for the need to employ more complex models that learn from user-item interactions, such as collaborative filtering (He et al., 2017). Finally, we think there is a lot of room for improvement in response generation, given the low performance of GPT-3.5 and our model in terms of BLEU and ROUGE, indicating that CPS is a promising direction but there is still a long way to go. To further validate the response generation, we add a case study shown in Figure 6. From the response generation example, we observe that, while GPT-3.5 generates general responses omitting product recommendations, our model generates reasonable responses with specific product recommendations, which are similar to the ground truth. However, the responses generated by our model do not contain the same phrase as the ground truth although their meaning is the same, leading to the relatively low BLEU and ROUGE.
@@ 6.2.3. Performance on System Action Prediction
The detailed performance of our model in terms of each system action, namely, ‘Chitchat’, ‘Clarify’, ‘Recommend’, and ‘No-answer’, is shown in Table 4. We observe that the performance changes for different system actions. Specifically, the model achieves the highest F1 when the system actions are ‘Recommend’ and ‘Clarify’. This might be because ‘Recommend’ and ‘Clarify’ actions are less ambiguous and can be easily classified.
@@ 7. Conclusion and Future Work
To collect the data, two workers need to be paired as participants to generate one conversation. Each conversation usually involves multiple searches and scanning for products, which costs a lot of effort for the participants. This makes the data collection challenging and expensive, leading to a limited volume of data like existing work (Trippas et al., 2020; Thomas et al., 2017; Dalton et al., 2020; Bernard and Balog, 2023). We understand that, although the data contains more than one thousand conversations, the data is still limited to training a large model in an end-to-end manner. To this end, we utilize other conversational datasets to pre-train our model to relax this limitation and the experimental results of our proposed model demonstrate its effectiveness. Nevertheless, we plan to expand the dataset with more training data from a large number of users in the future. As we present a system and data collection protocol in the paper, one can easily collect more conversations by using our system and protocol. Moreover, more than one thousand conversations enable researchers to use the dataset with LLMs (Mao et al., 2024) and unlock its capabilities, e.g., one can employ LLM’s few-shot learning by carefully curating domain-specific examples and prompts from the dataset; one can also use LLM’s knowledge injection by integrating our conversations as demonstrations or leveraging our knowledge graphs through retrieval-augmented generation (RAG) to enhance its reasoning and contextual understanding.
@@
In this paper, we experiment with our PSCon dataset to offer an overview of the difficulty and potential of our PSCon for targeting different tasks. Nevertheless, we do not experiment with cross-lingual or cross-market models (Zhu et al., 2022; Bonab et al., 2021) on our datasets. As we collect two datasets from two markets with two languages, there are several items from the international e-commerce stores that are the same across the two datasets. One can use the conversation in one market with a certain language to retrieve the target products in the other market with a different language. The information and knowledge of items can be transferred from one dataset to another dataset, which can be used for training domain-adaptation product search models. Also, the item interactions from conversations in one dataset can be learned and used as collaborative signals (e.g., users like Roccat Kain 120 AIMO mouse may also like Roccat Kone XP Air mouse, as stated in our PSCon dataset) in the other dataset for improving cross-domain (i.e., cross-lingual or cross-market) product recommendations. We plan to expand the dataset with more languages/markets and train cross-domain product search models (e.g., (Bonab et al., 2021) and (Zhu et al., 2022)) on our datasets in our future work.
Last, it is worth mentioning again that CPS is a promising direction and there is still a long way to go to improve the performance of each of the six subtasks.
Acknowledgements.
This research was supported by the National Natural Science Foundation of China (62402093) and, the Sichuan Science and Technology Program (2025ZNSFSC0479). This work was also supported in part by the National Natural Science Foundation of China under grants U20B2063 and 62220106008, and the Sichuan Science and Technology Program under Grant 2024NSFTD0034.